OpenStudy (anonymous):

Given the points S(-4,3) and T(-2,5), write an equation for which ST is the diameter.

OpenStudy (anonymous):

I got: (x+3)^2 + (y-4)^2 = 1

OpenStudy (anonymous):

am I correct?

OpenStudy (binary3i):

but not equal to 1

OpenStudy (binary3i):

\[=2\sqrt{2}\]

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

How'd you get that?

OpenStudy (binary3i):

sorry not that but equal to 8

OpenStudy (binary3i):

because that is the radius

OpenStudy (anonymous):

The general form of a circle with center \((h,k)\) is \((x-h)^2+(y-k)^2=r^2\)

OpenStudy (anonymous):

@binary3i I got √2 as the radius

OpenStudy (binary3i):

oh yeah you are correct.

OpenStudy (anonymous):

from what I got my answer is : (x+3)² + (y-4)² = 2

OpenStudy (anonymous):

(-4,3) isn't the center. Its a end point of the diameter.

OpenStudy (binary3i):

correct

OpenStudy (anonymous):

You are right! I need to read more carefully.

OpenStudy (anonymous):

so I went and found the midpoint and got (-3,4) and found the distance between the midpoint and one of the points given and got the square root of 2 I just want to know if my answer is right or if I made a mistake: (x+3)² + (y-4)² = 2

OpenStudy (anonymous):

You are right.

OpenStudy (anonymous):

Thank you very much :)